OpenStudy (anonymous):

Can someone explain why 1-3i/-2-4i simplies to 1/2 + 1/2i?

OpenStudy (anonymous):

just multiply by the complex conjugate (top and bottom) (-2+4i)/(-2+4i) and it works out :)

OpenStudy (anonymous):

can you supply the details?

OpenStudy (anonymous):

yeah, just multiply the top and bottom by (-2+4i) so expanding the top out you should get (1-3i)(-2+4i)= -2+4i+6i+12 On the bottom you'll get (-2-4i)(-2+4i)=4+16 You have (10+10i)/(20) which is your answer

OpenStudy (anonymous):

Thanks, but why is the i not squared?

OpenStudy (anonymous):

because I^2 =-1

OpenStudy (anonymous):

remember that i=sqrt(-1)

OpenStudy (anonymous):

I see. Thanks.